TAC Meeting Postmortem. DanaBalser summarized the discussion at the TAC postmortem.
Large Proposals. There were several subtopics. SRP Fatigue: Some SRPs see the same proposal each semester which could be submitted as a Large proposal. The TAC consensus was that there should not be strict rules. We briefly discussed issues with past Large proposals where the TAC would require multiple submissions; this practice is not longer an issue. Data Management (DM) Plan: The TAC suggested the DM plan be a separate document; 8 pages for the scientific justification and 2 pages for the DM plan. We discussed the need for a DM plan with the delivery of science ready data products. We still need a DM plan since the authors may want to produce their own data products and other components. It was not clear if this should include the observing plan. DanaBalser will check our current documentation. Large Proposal Reports: The TAC questioned if there should be more specific guidelines but felt they could deal with this on a case-by-case basis; they felt we should deliver the reports 2 weeks before the Large proposal meeting. We discussed if these reports should be more targeted, but the intent of these reports is a progress report (for the TAC and the observatory). We agreed that these should be availble well before and discussed at the Large proposal meeting; add to the agenda.
TAC schedule. The TAC discussed whether to spread the TAC meeting over more shorter days given remote format. There was no consensus by the TAC. MarkClaussen noted that this would give us less time to send out disposition letters.
Directors' Review. The VLA and VLBA materials have been submitted; still waiting for GBT and metrics. ToneyMinter and AnandCrossley said they should submit the info sometime today. ToneyMinter noted that the available hours for the GBT have not been reported correctly since the telescope painting has been trakced. This causes A semester available hours to be too high by about 300 hours and B semester available hours to be too high by about 50 hours. This may explain the GBT fluxuations detected in some of the metrics by the User's Committee. We need to flesh this out by the next User's Committee meeting.
PST Development. LorantSjouwerman reported that access to webtest is restored and we should be on track. RickLively provided some details. We should be able to start developement next week as planned but the development environment is not ideal. Developers cannot use a local sandbox for the database but must use webtest. This could cause some collisions with testers on webtest for the developers. Too, the connection to the webtest database is slow and moving forward we need a better solution (e.g., another environment). MarkClaussen noted that this may cause problems when we copy production to webtest. This usually does not occur during development.
Proposal Metric Issues. DanaBalser reported that we have run out of time trying to vet the metrics before the Directors' Review. We will pick this up mid-November. For 21A we should be able to use the GBT web service to update the PST. ToneyMinter asked about proposals with C- or those on Hold. The PST cannot accommodate these designations. So C- will be represented by C and propsoals on hold will have to be updated later.
Historical Large/Triggered VLBA Proposals. MarkClaussen reported that he has made some progress but has not updated the web pages yet.